# Survey on Modeling of Articulated Objects

### v1

@@ 4.2.2 Strategy and Supervision Signal
[XWY*09, ] 3 handcrafted -
[MYY*10, ] 3 handcrafted -
[SHL*14, ] 1 handcrafted -
[LWL*16, ] 3 handcrafted -
[HLV*17, ] 1 supervised distance func.
[YHL*18, ] 3 supervised displacement
[WZS*19, ] 1 supervised joint params.
RPM-Net [YHY*19] arbitrary supervised displacement
[ATK19, ] 1 supervised joint params.
[LWY*20, ] 1 supervised joint params.
[SCZ21, ] 1 self-supervised point corr.
ScrewNet [JLCN21] 1 supervised joint params.
[AFA*22, ] 1 supervised joint params.
[QJR*22, ] 1 supervised joint params.
OPD [JMSC22] 1 supervised joint params.
OPDMulti [SJSC23] 1 supervised joint params.
[LXX*22, ] 1 supervised joint params.
[LZH*23, ] 1 self-supervised point corr.
GAPartNet [GXZ*23] 1 supervised NPCS maps
[LSH*23, ] 1 supervised joint params.
@@
A-SDF [MQK*21] 1 supervised joint params.
[WCM*22, ] 1 supervised joint params.
Ditto [JHZ22] 1 supervised joint params.
CLA-NeRF [TLYS22] 1 handcrafted -
CARTO [HIZ*23] 1 supervised joint params.
PARIS [LMS23] 1 self-supervised multi-view RGB
NAP [LDS*23] 1 generative -
CAGE [LTMS23] 1 generative -
@@ 5.1 Representations of Articulated Motion
Deformation flow. Deformation flow is a general form of motion representation that can describe not only the rigid motion but also the motion as free-form deformation. The deformation flow describes the motion as a vector field, where each vector denotes the spatial displacement for each point in space. [YHL*18, ] and RPM-Net [YHY*19] both choose the deformation flow to represent the articulation motion as the output of their networks. They both leverage recurrent neural networks, such as LSTM, to directly predict the deformation flow from the input point cloud. An example of the deformation field prediction network is illustrated in fig. 9. And the training is supervised by the ground truth deformation flow. One of the advantages of using deformation flow is that it can handle the motion of the parts in a free-form manner. This flexibility is useful when the number of parts is unknown as the deformation on each point can be later grouped to form articulated parts in arbitrary numbers. The deformation flow can be particularly useful when modeling the non-trivial motion. For example, RPM-Net can model the opening and closing motion of the cover of an umbrella, which is a non-linear motion that cannot be represented by a simple rigid transformation. Similarly, A-SDF [MQK*21] and [WCM*22, ] also leverage the deformation field to represent articulation for the objects in the task of articulated object creation. They both leverage the neural field as a coordinate-based network to implicitly represent the deformation by conditioning on the articulation state of the object. Once the network is trained, the reconstructed object can be deformed to different states by querying the neural field. However, one of the limitations of using deformation flow or field is that it is less constrained than a strict form of rigid transformation when we are dealing with rigidly moving parts. To alleviate this issue, [YHL*18, ] tried to optimize the output deformation flow with an As-rigid-as-possible (ARAP) objective [SA07] to preserve the local rigidity of the parts. RPM-Net [YHY*19] also tried to enforce rigidity by adding a regularization term to the training process.
Joint parameters. The most common way to represent the motion for the articulated part is to use the joint parameters. This representation is standard in the robotics community, where the joint parameters can be directly recorded in the URDF file for simulation and control. The complete set of joint parameters includes the joint type, the joint axis, the joint state, and joint limits. The joint types that are most commonly seen in passively articulated objects are the 1 DoF joints, such as the revolute joint and the prismatic joint. The joint with more than 1 DoF, such as the ball joint, can also be seen in more complex objects, e.g. mechanical parts. An overview of the joint types that are covered in the literature is shown in fig. 1. The joint axis is represented by a direction vector as the axis orientation and a position vector as the location of the joint pivot. The joint state denotes the current articulation state of the joint, which is usually represented by the rotation angle for the revolute joint and the translational distance for the prismatic joint. The joint limits denote the range of motion that is allowed for the joint at two ends. An illustration of the joint parameters estimated from the input mesh is shown in fig. 8. Most of the existing work formulates the mobility analysis as a joint parameter estimation problem from various inputs, which can be converted into a combination of classification and regression problems. Usually, only a subset of these parameters are estimated as the output, depending on the specific task. While predicting the joint type with its axis, some works focus on estimating the current joint state of the object being observed [ATK19, LWY*20, LXX*22, LZH*23, GXZ*23] whose task is also termed as the part pose estimation problem. And other works aim to understand the joint limit [XWY*09, MYY*10, HLV*17, LWY*20, SCZ21], which can be more challenging as it requires the network to understand the physical constraints of the object.
@@
Kinematic tree. The kinematic tree serves as a hierarchical representation of the joints and parts of an articulated object. This structure is particularly crucial for depicting objects with multiple parts and joints, as it clearly outlines the dependencies between them An example of the kinematic tree is shown in fig. 3. Extracting the kinematic tree is a challenging problem since the kinematic structure varies from object to object with different numbers of parts and is structured in different topologies. Some existing works bypass this step by assuming the kinematic tree is predefined by the semantic object class [ATK19, LWY*20, LZH*23, MQK*21, WCM*22] or by simplifying the problem to a single part with a single joint [HLV*17, JLCN21, JHZ22, TLYS22, HIZ*23, LMS23]. And a large portion of the works dealing with multiple parts and joints also do not explicitly represent the kinematic tree in their model [YHL*18, WZS*19, YHY*19, SCZ21, SJSC23, GXZ*23]. There are only a few works that explicitly model the kinematic tree. The early work by [MYY*10, ] and [SHL*14, ] extract the kinematic tree from the input mesh. They share a similar idea to leverage the intersection of the part surfaces to identify the connection between the parts. [MYY*10, ] detected the contact area between the parts to determine the topology of the interaction graph, and [SHL*14, ] applied a segment-cluster process to compute the support-tree hierarchy from decomposed parts. Later works by [AFA*22, ] and [LXX*22, ] attempt to predict the kinematic tree from the input point cloud. [AFA*22, ] use a graph neural network to predict the labeling for each segmented part. [LXX*22, ] leverage an encoder-decoder architecture backboned by PointNet++ [QYSG17] to predict the connectivity between each pair of parts.
Figure 8: Examples of joint parameters output from [SHL*14, ] by decomposing an object into a mobility tree and estimating articulation with heuristic rules. Figure reproduced from original paper [SHL*14].
@@ 5.2 Assumption on the Articulation Model
The number of degrees of freedom (# DoF). Most of the existing works assume to only deal with the 1 DoF joints, including the revolute joint, the prismatic joint, and the helical joint. These joints are the most common types of joints that are seen in the daily articulated objects, such as cabinets, dishwashers, refrigerators, etc. And their motion is constrained to be linear and easily represented by a combination of rotation and translation constrained by a single joint axis. When dealing with more complex objects, such as delicate mechanical systems, the joint with more than 1 DoF, such as the ball joint, can also be seen. A few methods [XWY*09, MYY*10, LWL*16, YHL*18] have been proposed to deal with joint up to 3 DoF, which is achieved by either leveraging the deformation field to represent the motion or by combining with the handcrafted features on the part geometry to classify the joint type.
@@
Known object category (Obj. cat.) As mentioned above, one of the main challenges in articulation modeling is the diversity of the kinematic structure of the objects. One way to generalize the knowledge learned from the training data to the unseen objects is to build a category-level model, which is based on the assumption that the objects in the same category share the same structures. The structure implied by this line of work [ATK19, LWY*20, LZH*23] is not only about the kinematic structure but also the geometric structure. It means that the number of parts, how the parts are arranged spatially, and the way each part is articulated are all shared among the objects in the same category. This presents a strong prior for the model to learn from. In some cases, this level of category can be aligned with the semantic class of the objects whose kinematic structure is relatively simple and fixed, such as eyeglasses, laptops, scissors, etc. However, this assumption is highly constrained since this is not the case for more objects with more complex structures. For example, a refrigerator with two doors stacked vertically and a refrigerator with two doors stacked horizontally are considered to be from the same semantic class but with different structures. A separate model should be trained for each kind of refrigerator under this assumption. It makes the model less scalable to real-world applications with a diversity of objects.
@@
Known number of joints (# Joints). Whether the number of joints is known is another assumption made by the existing works. The previous assumption of knowing the object category inherently implies that the number of joints is known. There are also a few works [HLV*17, JLCN21, JHZ22, TLYS22, HIZ*23, LMS23] assume to estimate only one joint at a time but not knowing the motion the joint exhibits. It can be useful to generalize the method across objects with varied kinematic ways for simple objects. However, it is not scalable to more complex objects with more than one joint, highlighting a limitation in this line of work.
Figure 9: An illustration of using a recurrent neural network to predict deformation field for each articulated part from RPMNet [YHY*19]. Figure reproduced from original paper [YHY*19].
@@ 5.3 Methodologies for Articulation Modeling
Handcrafted methods. Early works leverage non-learning-based methods to infer the articulation model. [XWY*09, ] applied slippage analysis to compute a set of joint parameters from a mesh input. The output includes the number of rotational and translational degrees of freedom which is thresholded to be up to three, and the corresponding axes for each slippable motion. To determine the motion range for each joint, they design a trial-and-error bisection process to decide the feasibility. They estimate the joint limit by probing the possible motions until the motion results in an unrealistic geometry, e.g. penetration occurs between the parts. [MYY*10, ] defines a list of heuristic rules to determine the joint type, axis, and limit for each mechanical part. Once detecting the supporting-supported relationship between parts, [SHL*14, ] extract this connectivity to form a kinematic tree. By leveraging the reoccurrence of the objects and parts in the scene with different poses, they design rules to determine the motion type and axes by using the PCA axes of the part geometry. Taking a point cloud sequence as input, [LWL*16, ] propose a random sampling consensus method in 4D to fit motion on 3D trajectories. Then these motions can be clustered to convert to the joint parameters mathematically and the joints can be organized in a mobility graph.
Supervised learning. Most of the later works leverage supervised learning to predict the articulation model as shown in table 3. Depending on the articulation representation chosen, the training process can be supervised by the ground truth joint parameters or deformation flows. Different from these works that explicitly articulation annotation, [HLV*17, ] propose to classify the motion type from a mesh input through a metric learning model. The training is supervised by a distance function with several constraints to pull the object pairs with the same motion type closer and push the ones with different motion types away. Illustrated in fig. 10, [LWY*20, ] proposed a method to canonicalize an articulated object using coordinate spaces at object-level as Normalized Articulated Object Coordinate Space (NAOCS) and in part-level as Normalized Part Coordinate Space (NPCS). The canonical space is defined for each object category. Once the object is canonicalized, the joint parameters can be estimated by fitting the attributes on the template object. GAPartNet [GXZ*23] follows the definition of the above coordinate space for each part to supervise the training with the ground truth NPCS maps. After learning the NPCS mapping from the input state to a canonical state, they apply the Umeyama algorithm [Ume91] to estimate the rigid transformation from the mapping to obtain the joint parameters.
@@
Figure 10: Overview of the ANCSH method proposed by [LWY*20, ] that maps the object at an arbitrary articulated state to a hierarchical normalized coordinate space. Figure reproduced from original paper[LWY*20].
@@
Self-supervised learning. Although supervised learning is effective in learning the articulation model, it requires a large amount of labeled data to train the model. To alleviate this issue, a few works propose to leverage self-supervised learning to estimate the articulation model. [SCZ21, ] leverage a network to learn the point correspondence across different frames of the point cloud sequence from the input. Once the 3D trajectories are constructed by learning the consistent point correspondence, the joint parameters can be estimated by fitting the motion on the trajectories. One limitation of this work is that it requires observing a motion sequence as input, which usually involves human intervention while capturing the data. This tedious motion-capturing process makes the method less scalable to real-world applications. To alleviate this issue, [LZH*23, ] propose to learn the joint parameters from a single point cloud in a self-supervised manner. By factorizing each articulated part to its own normalized canonical space and then reconstructing it back to the input state using an optimized transformation, the reconstruction consistency in this process can be used to supervise the training. However, this work is limited by the assumption that the model is category-specific, meaning the geometric and kinematic structure of the object is known in advance. In the task of articulated object reconstruction, PARIS [LMS23] proposes to learn the joint parameters from a pair of multi-view RGB images self-supervisedly. The estimation of the joint parameters is supervised by the consistency of the photometric loss across two articulated states given as input. Although this method can be generalizable across object categories and does not rely on any 3D or articulation supervision, it assumes only one part is moving in the input observation.
6 Future Discussion and Conclusion
As the field of 3D modeling for articulated objects continues to evolve, several promising research directions have emerged from our survey. The key areas for future exploration include:

### v3

@@ 5 Articulation Modeling
xu2009joint 3 handcrafted -
mitra2010illustrating 3 handcrafted -
sharf2014mobility 1 handcrafted -
li2016mobility 3 handcrafted -
hu2017learning 1 SL distance func.
yi2018deep 3 SL displacement
wang2019shape2motion 1 SL joint params.
RPM-Net yan2019rpm arbitrary SL displacement
abbatematteo2019learning 1 SL joint params.
ANCSH li2020category 1 SL joint params.
CAPTRA weng2021captra 1 SL joint params.
shi2021self 1 SSL point corr.
ScrewNet jain2021screwnet 1 SL joint params.
MultiBodySync huang2021multibodysync 1 SSL point corr.
abdul2022learning 1 SL joint params.
qian2022understanding 1 SL joint params.
OPD jiang2022opd 1 SL joint params.
OPDMulti sun2023opdmulti 1 SL joint params.
liu2022toward 1 SL joint params.
liu2023self 1 SSL point corr.
liu2023building 1 SSL point corr.
GAPartNet geng2023gapartnet 1 SL NPCS maps
liu2023semi 1 SL joint params.
AutoURDF lin2024autourdf 1 SSL point corr.
GAMMA yu2024gamma 1 SL joint params.
@@
A-SDF mu2021asdf 1 SL joint params.
wei2022self 1 SL joint params.
Ditto jiang2022ditto 1 SL joint params.
CLA-NeRF tseng2022cla 1 handcrafted -
CARTO heppert2023carto 1 SL joint params.
liu2023few 1 TL, handcrafted -
PARIS liu2023paris 1 SSL MV RGBs
SfA nie2023sfa 1 SL labeled PC
NAP lei2023nap 1 generative joint params.
CAGE liu2024cage 1 generative joint params.
weng2024neural 1 SSL MV RGB-Ds
REACTO song2024reacto arbitrary SL optical flow
Real2Code mandi2024real2code 1 TL joint params.
RSRD kerr2024robot 1 SSL MV RGBs, video
SINGAPO liu2024singapo 1 generative joint params.
Articulate-Anything le2024articulate 1 TL -
MeshArt gao2024meshart 1 generative joint params.
ArtFormer su2024artformer 1 generative joint params.
Articulate AnyMesh qiu2025anymesh 1 TL, handcrafted -
ArtGS liu2025building 1 SSL MV RGB-Ds
@@ 5.1 Representations of Articulated Motion
Joint parameters. The most common way to represent part articulation is through joint parameters, which define the motion properties of articulated parts in terms of joint type, axis, state, and limits. This representation is widely used in robotics and simulation, as it provides a structured way to model how parts move relative to each other. The joint types most commonly seen are the 1 DoF joints, such as the revolute joint and the prismatic joint. The joint with more than 1 DoF, such as the ball joint, can also be seen in more complex objects, e.g. mechanical parts. An overview of the joint types that are covered in the literature is shown in Figure 1. The joint axis consists of a direction vector for orientation and a position vector for the pivot. The joint state captures its current configuration, typically as a rotation angle (revolute) or translation distance (prismatic). Joint limits define the allowed motion range to ensure physical plausibility.
The line of work on mobility analysis formulates joint parameter estimation as a combination of classification and regression problems, depending on the task requirements. Typically, joint type, position, and orientation are the primary components to be estimated. Some works additionally estimate the joint limit xu2009joint; mitra2010illustrating; hu2017learning; li2020category; shi2021self; qiu2025anymesh, which can be more challenging as it requires the network to understand the physical constraints of the object. Other works focus on estimating the pose of the articulated parts from 3D observation of objects abbatematteo2019learning; li2020category; weng2021captra; liu2022toward; liu2023self; geng2023gapartnet. Part pose is typically represented as a 3D amodal bounding box with a free 6D pose. By mapping the estimated pose to a normalized coordinate space introduced by li2020category, each part pose can be further converted into the joint state.
@@
Kinematic tree. The kinematic tree serves as a hierarchical representation of the joints and parts of an articulated object. This structure is particularly crucial for depicting objects with multiple parts and joints, as it clearly outlines the dependencies between them An example of the kinematic tree is shown in Figure 3. Extracting the kinematic tree is a challenging problem since the kinematic structure varies from object to object with different numbers of parts and is structured in different topologies. Some existing works bypass this step by assuming the kinematic tree is predefined per object class abbatematteo2019learning; li2020category; liu2023self; mu2021asdf; wei2022self, or by estimating only one joint at a time hu2017learning; jain2021screwnet; jiang2022ditto; tseng2022cla; heppert2023carto; liu2023paris. And a large portion of the works dealing with multiple joints also do not represent the kinematic tree in their model yi2018deep; wang2019shape2motion; yan2019rpm; shi2021self; sun2023opdmulti; geng2023gapartnet as they assume that all the articulated parts move relative to the static part.
There are only a few works that explicitly model the kinematic tree as the output of the mobility analysis. The early works by mitra2010illustrating and sharf2014mobility extract the kinematic tree from the input mesh. They share a similar idea to leverage the intersection of the part surfaces to identify the connection between the parts. mitra2010illustrating detected the contact area between the parts to determine the topology of the interaction graph, and sharf2014mobility applied a segment-cluster process to compute the support-tree hierarchy from decomposed parts. Later works by abdul2022learning and liu2022toward attempt to predict the kinematic tree from the input point cloud by labeling the dependency existence between each pair of segmented parts. Articulate AnyMesh qiu2025anymesh predicts the kinematic dependency using VLMs. For the work focusing on articulated object generation lei2023nap; liu2024cage; liu2024singapo; le2024articulate; gao2024meshart; su2024artformer, the kinematic tree is an essential output, as their ultimate goal is to synthesize fully articulatable assets that can be directly integrated into the simulation environment.
Deformation flow. Deformation flow is a general form of motion representation that can describe not only the rigid motion but also the motion as free-form deformation, making it highly versatile for modeling articulation. It describes the motion as a vector field, where each vector denotes the spatial displacement for each point in space. Unlike traditional joint-based representations, deformation flow allows for more flexible motion modeling, especially when the motion involves non-trivial transformations.
For mobility analysis, yi2018deep and RPM-Net yan2019rpm opt to use deformation flow to represent the articulation motion. They both leverage recurrent neural networks, such as LSTM, to directly predict the deformation flow from the input point cloud. An example of the deformation field prediction network is illustrated in Figure 13. One of the advantages of using deformation flow is that it can handle the motion of the parts in a free-form manner. This flexibility is useful when the number of parts is unknown as the deformation on each point can be later grouped to form articulated parts in arbitrary numbers. The deformation flow can be particularly useful when modeling the non-trivial motion. For example, RPM-Net can model the opening and closing motion of the cover of an umbrella, which is a non-linear motion that cannot be represented by a simple rigid transformation.
Similarly, A-SDF mu2021asdf, wei2022self, and REACTO song2024reacto also leverage the deformation field to represent articulation for the objects in the task of articulated object creation. These methods optimize coordinate-based networks to implicitly represent the deformation, conditioned on the articulation state of the object. Once trained, the network allows the reconstructed object to be deformed into different states by querying the neural field at different timesteps, enabling continuous articulation modeling.
However, a key limitation of deformation flow or fields is that they are less constrained than strict rigid transformations, which can be problematic when dealing with rigidly moving parts. To alleviate this issue, yi2018deep proposes to optimize the deformation flow with an as-rigid-as-possible (ARAP) objective sorkine2007rigid to preserve the local rigidity of the parts. RSRD kerr2024robot estimates articulation by tracking the pose for each cluster of 3D Gaussians over time in an input video. By treating 3D Gaussians as soft points in space, an ARAP loss penalizes deviation between neighboring Gaussians, ensuring that deformation follows locally rigid motion. This approach bridges deformation-based articulation modeling with Gaussian-based representation learning, demonstrating the potential for more flexible yet structured articulation tracking.
Discussion on articulation representations. Joint parameters, kinematic trees, and deformation fields are three common ways to represent articulation, each with its own strengths and challenges. Joint parameters are fundamental for describing discrete motion constraints and are widely used in robotics and simulation. Kinematic trees define the hierarchical structure of articulation dependency, which complements joint parameters to form a complete articulation model while also serving as a crucial component for simulation-ready articulated assets. While often overlooked due to the simplicity of many articulated objects with single-level structures, kinematic trees become essential when handling objects with deeper articulation hierarchy (e.g., table lamps with multiple arms, robotic mechanisms, mechanical assemblies). Modeling such structures introduces additional challenges for mobility analysis and articulated asset creation, as it requires reasoning over highly structured data. In contrast, deformation fields offer a continuous and flexible representation, capable of capturing both rigid and non-rigid motion. However, they lack explicit structural constraints when modeling piece-wise rigidity, usually requiring additional regularization techniques to enforce local rigidity.
Figure 13: An illustration of using a recurrent neural network to predict deformation field for each articulated part from RPMNet yan2019rpm. Figure reproduced from original paper yan2019rpm.
@@ 5.2 Assumption on the Articulation Model
The number of degrees of freedom (# DoF). Most of the existing works assume to only deal with the 1 DoF joints, including the revolute joint, the prismatic joint, and the helical joint. These joints are the most common types of joints that are seen in the daily articulated objects, such as cabinets, dishwashers, refrigerators, etc. And their motion is constrained to be linear and easily represented by a combination of rotation and translation constrained by a single joint axis. When dealing with more complex objects, such as delicate mechanical systems, the joint with more than 1 DoF, such as the ball joint, can also be seen. A few methods xu2009joint; mitra2010illustrating; li2016mobility; yi2018deep have been proposed to deal with joints up to 3 DoF, which is achieved by either leveraging the deformation field to represent the motion or by combining with the handcrafted features on the part geometry to classify the joint type.
@@
Known object category (Obj. cat.) As mentioned above, one of the main challenges in articulation modeling is the diversity of the kinematic structure of the objects. One way to generalize the knowledge learned from the training data to the unseen objects is to build a category-level model, which is based on the assumption that the objects in the same category share the same structures. The structure implied by this line of work abbatematteo2019learning; li2020category; liu2023self is not only about the kinematic structure but also the geometric structure. It means that the number of parts, how the parts are arranged spatially, and the way each part is articulated are all shared among the objects in the same category. This presents a strong prior for the model to learn from. In some cases, this level of category can be aligned with the semantic class of the objects whose kinematic structure is relatively simple and fixed, such as eyeglasses, laptops, scissors, etc. However, this assumption is highly constrained since this is not the case for more objects with more complex structures. For example, a refrigerator with two doors stacked vertically and a refrigerator with two doors stacked horizontally are considered to be from the same semantic class but with different structures. A separate model should be trained for each kind of refrigerator under this assumption. It makes the model less scalable to real-world applications with a diversity of objects.
@@
Known number of joints (# Joints). Whether the number of joints is known is another assumption made by the existing works. The previous assumption of knowing the object category inherently implies that the number of joints is known. There are also a few works hu2017learning; jain2021screwnet; jiang2022ditto; tseng2022cla; heppert2023carto; liu2023paris that assume to estimate only one joint at a time but not knowing the motion the joint exhibits. It can be useful to generalize the method across objects with varied kinematic ways for simple objects. However, it is not scalable to more complex objects with more than one joint, highlighting a limitation in this line of work.
@@ 5.3 Methodologies for Articulation Modeling
Handcrafted methods. Early works leverage non-learning-based methods to infer the articulation model. xu2009joint applied slippage analysis to compute a set of joint parameters from a mesh input. The output includes the number of rotational and translational degrees of freedom which is thresholded to be up to three, and the corresponding axes for each slippable motion. To determine the motion range for each joint, they design a trial-and-error bisection process to decide the feasibility. They estimate the joint limit by probing the possible motions until the motion results in an unrealistic geometry, e.g. penetration occurs between the parts. mitra2010illustrating define a list of heuristic rules to determine the joint type, axis, and limit for each mechanical part. Once detecting the supporting-supported relationship between parts, sharf2014mobility extract this connectivity to form a kinematic tree. By leveraging the reoccurrence of the objects and parts in the scene with different poses, they design rules to determine the motion type and axes by using the PCA axes of the part geometry. Taking a point cloud sequence as input, li2016mobility propose a random sampling consensus method in 4D to fit motion on 3D trajectories. Then these motions can be clustered to convert to the joint parameters mathematically and the joints can be organized in a mobility graph. CLA-NeRF tseng2022cla assumes the object has a single revolute joint whose axis is at the intersection of the two articulated segments. So once the parts are segmented in the NeRF, the joint axis can be directly voted by the points near the intersected area by applying linear regression.
@@
Figure 14: Overview of the ANCSH method proposed by li2020category that maps the object at an arbitrary articulated state to a hierarchical normalized coordinate space. Figure reproduced from original paperli2020category.
Supervised learning. A recent line of work focusing on mobility analysis leverages supervised learning to predict the articulation model from a single state of the object. Depending on the articulation representation chosen and perception input, the training process can be supervised by the ground truth motion parameters wang2019shape2motion; jain2021screwnet; qian2022understanding; nie2023sfa or point displacements yi2018deep; yan2019rpm. As a representative work shown in Figure 8, Shape2Motion wang2019shape2motion designs a motion attribute proposal module to decode point-wise features into parameters for all the possible articulation in the object. Rather than directly regressing the joint axis, they reformulate joint location prediction as a point selection problem and the joint orientation problem as a classification task over 14 discrete directions, which reduces the complexity of direct regression in high-dimensional continuous spaces. Once each motion is proposed, a matching module follows to pair articulations with the corresponding parts, allowing for cases where a single articulated part may exhibit multiple 1-DoF motions. Different from the above works that explicitly use motion annotation, hu2017learning propose to classify the motion type from a mesh input through a metric learning model. The training is supervised by a distance function with several constraints to pull the object pairs with the same motion type closer and push the ones with different motion types away.
Another line of work tackles the task of articulated part pose estimation from the depth image of an object in an arbitrary articulation state in a supervised manner. By assuming the articulation structure is shared within each object category, li2020category pioneers in designing an articulation-aware normalized coordinate space hierarchy (ANCSH), which is used to canonicalize the perceived object at the part level. As shown in Figure 9, ANCSH is composed of a Normalized Articulated Object Coordinate Space (NAOCS) on top of a set of Normalized Part Coordinate Spaces (NPCSs) per part, and the joint parameters in the resting state have been pre-defined in NAOCS. Once this hierarchical representation is defined for each object category, a deep neural network (illustrated in Figure 14) is trained to predict the mapping from each point to a coordinate in NPCS, followed by a head to predict the transformation from each NPCS to NAOCS and another head to predict the point-joint association and joint parameters in NAOCS. The training is supervised by multiple losses using ground truth outputs, including per-point NPCS and NAOCS mapping, per-part transformations, joint parameters, etc. Leveraging this ANCSH representation, CAPTRA weng2021captra and GAPartNet geng2023gapartnet extend the problem setting to pose tracking and domain-generalizable object perception.
@@
Self-supervised learning. Although supervised learning is effective in learning the articulation model, a large amount of labeled data is required to train the model. To alleviate this issue, a few works shi2021self; huang2021multibodysync; liu2023self; lin2024autourdf propose to leverage self-supervised learning to estimate the articulation model by leveraging the correspondence across multi-state observations of the object. shi2021self leverage a network to learn the point correspondence across different frames of the point cloud sequence from the input. Once the 3D trajectories are constructed by learning the consistent point correspondence, the joint parameters can be estimated by fitting the motion on the trajectories. One limitation of this work is that it requires observing a motion sequence as input, which usually involves human intervention while capturing the data. This tedious motion-capturing process makes the method less scalable to real-world applications. To alleviate this issue, liu2023self propose to learn the joint parameters from a single point cloud in a self-supervised manner. By factorizing each articulated part to its own normalized canonical space and then reconstructing it back to the input state using an optimized transformation, the reconstruction consistency in this process can be used to supervise the training. However, this work is limited by the assumption that the model is category-specific, meaning the geometric and kinematic structure of the object is known in advance.
Figure 15: An illustration of a self-supervised learning method proposed in PARIS liu2023paris that uses photometric consistency to jointly optimize part geometry and joint parameters . Figure reproduced from original paper liu2023paris.
In the task of articulated object reconstruction, PARIS liu2023paris proposes to learn the joint parameters from a pair of multi-view RGB images in a self-supervised manner. As illustrated in Figure 15, the estimation of the joint parameters is supervised by the consistency of a photometric loss across two articulated states given as input. Although this method can be generalizable across object categories and does not rely on any 3D or articulation supervision, it assumes only one part is moving in the input observation. weng2024neural and ArtGS liu2025building follows a similar idea to optimize the joint parameters from a pair of observations in multi-view RGB-D images. Reconstruction losses on the RGB-D frames are used to first segment the parts in two states of the object, and then the joint parameters can be computed by optimizing the point correspondences between the segments from the two states. RSRD kerr2024robot models the object dynamics from human demonstration videos using a photometric consistency loss and regularization terms to constrain the motion of the parts with as-rigid-as-possible (ARAP) sorkine2007rigid penalty. This work tracks the articulated part in 4D poses along the time steps, so it can be transferred to a physical grasping robot to execute these motions.
Generative methods. As mentioned in Section 4.2.2, there is a line of work lei2023nap; liu2024cage; liu2024singapo; gao2024meshart; su2024artformer that trains generative models from scratch to learn the joint distribution of geometry and articulation for each part. On the articulation side, all existing methods model the motion properties explicitly through joint parameters. The most common approach liu2024cage; liu2024singapo; gao2024meshart; su2024artformer is to parameterize joints with joint location, orientation, and motion limits, with some methods also including joint type liu2024cage; liu2024singapo and a joint existence indicator lei2023nap; gao2024meshart. NAP lei2023nap adopts Plucker coordinates for joint representation, using a unit direction vector $l$ and momentum $m$ perpendicular to $l$, providing a unified representation for prismatic and revolute joints. Notably, all the above joint parameters are defined in the global object frame, which is consistent with the part shape representation. During training, the loss function is designed to enforce consistency with the ground-truth joint parameters per part.
However, an open question remains as to whether a global-frame or local-frame parameterization is more efficient in the generative context. The global-frame representation captures the distribution across different objects in a coordinated manner, but it is sensitive to the object’s orientation and articulation states. A more local representation can potentially capture the connection within the part geometry in a more compact way. Exploring alternative formulations to compactly encode joint parameters alongside part geometry is a promising direction for future research, leading to more scalable and generalizable generative models for articulated objects.
Figure 16: The actor-critic program system for link placement and joint prediction proposed in Articulate-Anything le2024articulate. Figure reproduced from the original paper le2024articulate.
Transfer learning methods. Real2Code mandi2024real2code and Articulate-Anything le2024articulate are two recent works that leverage pre-trained LLMs to transfer language reasoning and code generation capabilities to the articulation understanding problem. These methods bridge the gap between articulation modeling and generative programming, enabling articulated object representations that are both interpretable and simulation-ready. Real2Code abstracts the part layout as oriented bounding boxes after segmenting articulated parts. It then defines joint parameters relative to these bounding boxes, formulating joint prediction as a classification problem. Taking the bounding boxes as input, they fine-tune CodeLlama roziere2023codellama to predict the joint parameters and generate code to construct the objects into a compact representation that can be directly used in physics simulations. In contrast, Articulate-Anything builds a joint prediction system on VLMs. It takes code representations that configure link placements as input and generates kinematic parameters between parts. The prediction is iteratively refined by taking feedback from another program that criticizes the realism of the results by comparing it with the input image or video. Both approaches demonstrate the potential of LLMs and VLMs in articulation reasoning and code-based object generation. Articulate AnyMesh qiu2025anymesh also leverages VLMs to analyze the part articulation. Based on some heuristic rules, they prompt GPT-4o to make selections from several candidate markers on the rendered images or possible joint directions, depending on the joint type.
Figure 17: Example sources of human-in-the-loop video: human interactions, egocentric videos, and human demonstrations videos that involve articulated objects. The figures are reproduced from the original papers qian2022understanding; grauman2022ego4d; kerr2024robot.
6 Discussion and Conclusion
This survey paper provided an overview of the current state-of-the-art in 3D modeling of human-made articulated objects, focusing on the tasks of articulated part perception and articulated object creation (reconstruction and generation). We first defined the scope of the discussion and provided a compilation of articulated object datasets used in the research community. We then analyzed techniques and methodologies developed for 3D modeling of articulated objects, focusing on two key aspects: 1) geometry modeling — representation and approaches used to understand the shape of articulated parts and 2) articulation modeling — articulation models and methodologies used to estimate part mobility and kinematic structures. We highlighted gaps in the literature and opportunities for further work emerging from our survey, which are likely to be fruitful as the field of 3D modeling for articulated objects continues to evolve. We conclude by identifying key challenges and potential research directions for future exploration.